# Learning to Optimize for Mixed-Integer Non-linear Programming

### v4

@@ 3 Learning to Optimize MINLPs: a Problem Formulation
$\mathcal{L}(\Theta)=\frac{1}{m}\sum_{i=1}^{m}\Big{(}{\bf f}({\bf\hat{x}}^{i},{\bm{\xi}}^{i})+\lambda\cdot\left({\bf g}({\bf\hat{x}}^{i},{\bm{\xi}}^{i})\right)_{+}\Big{)}\text{ with }{\bf\hat{x}}^{i}=\bm{\psi}_{\Theta}({\bm{\xi}}^{i}),$ (2)
where $\left(\cdot\right)_{+}$ ensures only positive constraint violations are penalized (implemented via a ReLU function), and $\lambda>0$ is a penalty hyperparameter that balances the trade-off between minimizing the objective function and satisfying the constraints.
@@ Methods.
Table 1 provides an overview of all the methods used in the following experiments. A 60 or 1000-second time limit is enforced for all methods and problems. The experiments evaluate our learning-based methods, Rounding Classification (RC) and Learnable Threshold (LT), against traditional exact optimization (EX), which can compute optimal solutions but is often computationally expensive, and heuristic-based approaches such as Rounding after Relaxation (RR) and root node solutions (N1), which offer faster results without quality guarantees. Note that baselines EX and N1 include a wide range of heuristics that are embedded in the MINLP solver of choice (Gurobi or SCIP) and that are executed in conjunction with the tree search procedure; we are also implicitly comparing to these heuristics, not just to the exact search. As such, the competing methods cover a broad spectrum of optimization strategies, from exact solvers to fast heuristics, allowing for a comprehensive evaluation of solution quality and computational efficiency. In addition, we evaluate two ablation baselines, which isolate different aspects of our correction layers $\varphi_{\Theta_{2}}$ to highlight their impact on performance. Details of these ablation studies, including methodology and results, are provided in Appendix F.
@@
Figure 2: Illustration of objective value evolution for a $100\times 100$ Convex Quadratic instance and a $60\times 4$ Rosenbrock instance over 600 seconds.
@@ 6.2 Convex Quadratic Problem
RC Obj Mean $0.803$ $-1.476$ $-4.151$ $-12.422$ $-13.564$ $-30.799$ $-71.373$ $-120.378$
Obj Median $0.648$ $-1.719$ $-4.166$ $-12.416$ $-13.634$ $-30.789$ $-71.476$ $-120.492$
% Infeasible $0\%$ $3\%$ $4\%$ $0\%$ $4\%$ $1\%$ $4\%$ $3\%$
Time (Sec) $0.0019$ $0.0019$ $0.0019$ $0.0019$ $0.0021$ $0.0025$ $0.0026$ $0.0045$
LT Obj Mean $0.743$ $-1.615$ $-4.165$ $-12.355$ $-13.395$ $-29.728$ $-71.350$ $-114.434$
Obj Median $0.554$ $-1.928$ $-4.203$ $-12.352$ $-13.527$ $-29.808$ $-71.601$ $-114.522$
% Infeasible $1\%$ $1\%$ $0\%$ $4\%$ $3\%$ $0\%$ $3\%$ $1\%$
Time (Sec) $0.0019$ $0.0019$ $0.0019$ $0.0019$ $0.0023$ $0.0022$ $0.0026$ $0.0046$
@@
Since there is a lack of publicly available datasets for parametric MINLPs, the convex quadratic problems used in the experiments are adapted from Donti et al. (2021), which originally focused on learning under continuous constraints. We introduced integrality constraints on all decision variables to tailor these problems to our discrete setting. Additionally, we removed equality constraints to avoid the issue of generating infeasible instances. These modifications ensure compatibility with our framework while preserving the essential structure of the original problems:
where the coefficients $\bm{Q}\in\mathbb{R}^{n\times n}$, $\bm{p}\in\mathbb{R}^{n}$, and $\bm{A}\in\mathbb{R}^{m\times n}$ were fixed, while $\bm{b}\in\mathbb{R}^{m}$ were treated as parametric coefficients (input features), varying across instances. Further details on the generation of $\bm{Q},\bm{p},\bm{A},$ and $\bm{b}$ can be found in Appendix E.
@@ 6.3 Simple Non-convex Problem
To explore the performance on non-convex optimization tasks, Donti et al. (2021) extended the convex quadratic programming problem by adding a trigonometric term to the objective function. Similarly, we modified the formulation from the convex quadratic as follows:
where the sine function is applied element-wise to the decision variables $\bm{x}$. This introduces non-convexity into the problem, making it more challenging compared to the convex case. The constraint matrix $\bm{A}$ is parametrized by $\bm{d}$ (details in Appendix E), an additional variable that influences the matrix structure, further enriching the complexity. Alongside $\bm{d}$, the right-hand side vector $\bm{b}$ remains a dynamic parameter in the problem. The coefficients $\bm{Q}$, $\bm{p}$, $\bm{A}$, and $\bm{b}$ are generated in the same way as in the quadratic formulation. In addition, the scales of the problem and the experiment setting are also identical to those of the quadratic problems.
Table 3: Results for a Simple Non-convex Problem. See the caption of Table 2 for details. The “—” symbol indicates that no solution is found for any instance within 60 seconds.
@@
RC Obj Mean $0.383$ $1.102$ $0.228$ $0.771$ $1.664$ $1.472$ $0.526$ $1.422$
Obj Median $0.240$ $0.838$ $0.217$ $0.752$ $1.594$ $1.436$ $0.526$ $0.809$
% Infeasible $1\%$ $3\%$ $0\%$ $2\%$ $0\%$ $1\%$ $4\%$ $3\%$
Time (Sec) $0.0019$ $0.0019$ $0.0019$ $0.0020$ $0.0022$ $0.0022$ $0.0029$ $0.0040$
LT Obj Mean $0.359$ $0.910$ $0.195$ $0.580$ $0.669$ $-0.356$ $-1.374$ $-3.744$
Obj Median $0.226$ $0.683$ $0.175$ $0.566$ $0.649$ $-0.373$ $-1.594$ $-3.716$
% Infeasible $0\%$ $2\%$ $1\%$ $2\%$ $4\%$ $0\%$ $2\%$ $1\%$
Time (Sec) $0.0019$ $0.0019$ $0.0019$ $0.0020$ $0.0021$ $0.0023$ $0.0029$ $0.0050$
EX Obj Mean $0.001$ $-0.182$ $-0.406$ $25.750$
Obj Median $-0.095$ $-0.310$ $-0.420$ $9.520$
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$
Time (Sec) $0.118$ $2.509$ $60.153$ $60.125$
RR Obj Mean $-0.047$ $-0.168$ $-0.464$ $-1.039$ $-2.068$ $-3.990$ $-7.935$
Obj Median $-0.089$ $-0.325$ $-0.476$ $-1.215$ $-2.307$ $-4.327$ $-7.104$
% Infeasible $64\%$ $86\%$ $97\%$ $100\%$ $100\%$ $100\%$ $56\%$
Time (Sec) $0.216$ $0.422$ $1.013$ $1.198$ $4.654$ $52.242$ $62.006$
N1 Obj Mean $1.690$ $31.1e$ 3 $42.1e$ 4 $63.7e$ 6
Obj Median $0.183$ $0.557$ $2.222$ $45.847$
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$
Time (Sec) $0.0434$ $0.1029$ $0.1516$ $8.936$
@@ 6.4 Multi-Dimensional Mixed-Integer Rosenbrock Problem
The parametric, high-dimensional mixed-integer Rosenbrock problem serves as a challenging testbed. Adapted from the classic Rosenbrock function, this formulation introduces additional complexity by incorporating integer variables, non-linear constraints, and parametric variations in high dimensions. The problem can be expressed as follows:
where $\bm{x}\in\mathbb{R}^{n}$ are continuous decision variables and $\bm{y}\in\mathbb{Z}^{n}$ are integer decision variables. The vectors $\bm{p}\in\mathbb{R}^{n}$ and $\bm{q}\in\mathbb{R}^{n}$ are fixed for each instance, while the parameters $b$ and $\bm{a}$ vary. Details on how the parameters $\bm{p},\bm{q},\bm{a},$ and $b$ were generated can be found in Appendix E. Due to the structure of the constraints, it is straightforward to always construct a trivial feasible integer solution, ensuring that the problem remains feasible across instances.
@@
Table 4: Results for the Mixed-Integer Rosenbrock Problem. The number of decision variables varies from 2 to 20,000, while the number of constraints is 5. See the caption of Table 2 for details. The “—” symbol indicates that no solution is found for any instance within 60 sec time limits. “Unsolved” denotes the percentage of instances that could not be solved within the given time limit.
@@
RC Obj Mean / Median $23.27/21.48$ $59.39/48.86$ $503.51/461.71$ $5938.37/5792.52$ $66883.20/66797.11$
% Infeasible $3\%$ $0\%$ $1\%$ $1\%$ $24\%$
Time (Sec) $0.0019$ $0.0019$ $0.0021$ $0.0033$ $0.0116$
LT Obj Mean / Median $23.18/20.80$ $62.51/63.40$ $622.78/626.04$ $5611.78/5557.82$ $47622.18/34518.87$
% Infeasible $2\%$ $0\%$ $0\%$ $3\%$ $34\%$
Time (Sec) $0.0019$ $0.0020$ $0.0026$ $0.0030$ $0.0127$
EX Obj Mean / Median $19.62/18.20$ $65.50/59.16$ $911.909.93e5/$ 911.90 $9262.092.50e11/$ 9262.09
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$
% Unsolved $0\%$ $0\%$ $0\%$ $8\%$ $100\%$
Time (Sec) $3.5630$ $60.2459$ $60.1416$ $60.3500$
RR Obj Mean / Median $22.24/22.19$ $51.171.20e4/$ 51.17 $501.901.43e4/$ 501.90 $85346.617.02e8/$ 85346.61
% Infeasible $45\%$ $41\%$ $18\%$ $1\%$
% Unsolved $0\%$ $0\%$ $42\%$ $92\%$ $100\%$
Time (Sec) $0.1877$ $0.5635$ $1.4163$ $8.4646$
N1 Obj Mean / Median $40.37/27.93$ $87.83/77.34$ $957.423.72e8/$ 957.42 $9325.378.31e12/$ 9325.37
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$
% Unsolved $0\%$ $0\%$ $0\%$ $6\%$ $100\%$
Time (Sec) $0.0313$ $0.0829$ $0.2241$ $12.4214$
@@ 6.6 Effect of Training Sample Size
To assess the impact of sample size, we trained the model on datasets of 800, 8,000, and 80,000 instances, adjusting training epochs to 2000, 200, and 20 (with early stopping) to ensure comparable iterations. All other hyperparameters remained consistent to isolate the effect of sample size.
@@
As shown in Figure 4, increasing the sample size yields significant improvements in both objective values and feasibility. With 80,000 samples for training, the infeasibility ratio was reduced to 5% on the test set, demonstrating that larger sample sizes help the model generalize better to unseen instances, particularly in extremely high-dimensional settings. These results underscore the importance of adequate sample size when training models. This strategy is especially practical with our approach due to its self-supervised nature, which bypasses the need for labeled optimal solutions.
@@ 7 Conclusion
Our method allows for solving large-scale parametric MINLPs heuristically, leading to significantly better solutions and computational efficiency; however, our methods do not guarantee feasibility. Future work could focus on furthering the feasibility guarantees through alternative constraint handling or solution post-processing. For specific problem classes, it may be possible to relax only a subset of the constraints into the learning loss function while optimizing over the remaining constraints directly using existing differentiable optimization layers, e.g., from Agrawal et al. (2019). Another potential extension of our work is to redesign the neural network architecture to accommodate a varying number of instance parameters and decision variables. This can be achieved by the use of set-based, permutation-equivariant architectures including but not limited to graph neural networks (Cappart et al., 2023; Dumouchelle et al., 2024; Chen et al., 2022b; 2024).
@@ Appendix A Gumbel-Sigmoid Trick
The Gumbel-Sigmoid trick works by adding noise to the logits $h$, in which noise is sampled from a Gumbel distribution, defined as:
@@
The Gumbel-Sigmoid function produces a soft approximation with random perturbation:
@@
where $g_{1}$ and $g_{2}$ are independent samples from the Gumbel distribution, and the temperature parameter $\tau$ controls the smoothness: as $\tau\rightarrow 0$, the output becomes closer to a hard binarization. In our experiments, we set $\tau=1$ for simplicity.

### v5

@@ 3 Learning to Optimize MINLPs: a Problem Formulation
$\mathcal{L}(\Theta)=\frac{1}{m}\sum_{i=1}^{m}\Big{[}{\bf f}({\bf\hat{x}}^{i},{\bm{\xi}}^{i})+\lambda\cdot\|{\bf g}({\bf\hat{x}}^{i},{\bm{\xi}}^{i})_{+}\|_{1}\Big{]}\text{ with }{\bf\hat{x}}^{i}=\bm{\psi}_{\Theta}({\bm{\xi}}^{i}),$ (2)
where $\|\left(\cdot\right)_{+}\|_{1}$ ensures only the sum of positive constraint violations are penalized (implemented via a ReLU function), and $\lambda>0$ is a penalty hyperparameter that balances the trade-off between minimizing the objective function and satisfying the constraints.
@@ Methods.
Table 1 provides an overview of all the methods used in the following experiments. A 1000-second time limit is enforced for all methods and problems. The experiments evaluate our learning-based methods, Rounding Classification (RC) and Learnable Threshold (LT), against traditional exact optimization (EX), which can compute optimal solutions but is often computationally expensive, and heuristic-based approaches such as Rounding after Relaxation (RR) and root node solutions (N1), which offer faster results without quality guarantees. Note that baselines EX and N1 include a wide range of heuristics that are embedded in the MINLP solver of choice (Gurobi or SCIP) and that are executed in conjunction with the tree search procedure; we are also implicitly comparing to these heuristics, not just to the exact search. As such, the competing methods cover a broad spectrum of optimization strategies, from exact solvers to fast heuristics, allowing for a comprehensive evaluation of solution quality and computational efficiency. In addition, we evaluate two ablation baselines, which isolate different aspects of our correction layers $\varphi_{\Theta_{2}}$ to highlight their impact on performance. Details of these ablation studies, including methodology and results, are provided in Appendix F.
@@
Figure 2: Illustration of objective value evolution for a $100\times 100$ Convex Quadratic instance and $60\times 4$ Rosenbrock instance over 600 seconds.
@@ 6.2 Convex Quadratic Problem
RC Obj Mean $0.827$ $-0.773$ $-3.859$ $-11.967$ $-12.809$ $-29.830$ $-67.518$ $-129.088$
Obj Median $0.641$ $-1.180$ $-3.888$ $-12.013$ $-12.760$ $-29.965$ $-67.633$ $-129.263$
% Infeasible $0\%$ $1\%$ $0\%$ $1\%$ $1\%$ $2\%$ $0\%$ $0\%$
Time (Sec) $0.0019$ $0.0019$ $0.0021$ $0.0019$ $0.0021$ $0.0025$ $0.0026$ $0.0045$
LT Obj Mean $0.869$ $-1.419$ $-3.830$ $-11.606$ $-12.227$ $-30.688$ $-69.136$ $-120.480$
Obj Median $0.672$ $-1.702$ $-3.818$ $-11.736$ $-12.299$ $-31.030$ $-69.222$ $-120.651$
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $7\%$ $0\%$ $5\%$
Time (Sec) $0.0019$ $0.0019$ $0.0019$ $0.0019$ $0.0024$ $0.0026$ $0.0028$ $0.0047$
@@
Since there is a lack of publicly available datasets for parametric MINLPs, the convex quadratic problems used in the experiments are adapted from Donti et al. (2021), which originally focused on learning under continuous constraints. We introduced integrality constraints on all decision variables to tailor these problems to our discrete setting. Additionally, we removed equality constraints to avoid the issue of generating infeasible instances. These modifications ensure compatibility with our framework while preserving the essential structure of the original problems. Further details on the mathematical formulation and data generation process can be found in Appendix E.
@@ 6.3 Simple Non-convex Problem
To evaluate the performance on non-convex optimization tasks, we extended the convex quadratic programming problem by introducing a trigonometric term to the objective function, following the approach in Donti et al. (2021). This modification introduces non-convexity, increasing the challenge of finding optimal solutions. Additionally, we parameterized the constraint matrix to further enrich the complexity. Further details on the formulation, parameter generation, and experimental setup can be found in Appendix E. In addition, the scales of the problem and the experiment setting are also identical to those of the quadratic problems.
Table 3: Results for a Simple Non-convex Problem. See the caption of Table 2 for details. “Unsolved” denotes the percentage of instances that could not be solved within the given time limit.
@@
RC Obj Mean $0.855$ $1.236$ $0.598$ $1.352$ $1.208$ $1.522$ $-0.351$ $4.051$
Obj Median $0.496$ $0.865$ $0.601$ $1.325$ $1.198$ $1.522$ $-0.341$ $3.869$
% Infeasible $1\%$ $0\%$ $0\%$ $0\%$ $1\%$ $0\%$ $1\%$ $0\%$
Time (Sec) $0.0019$ $0.0019$ $0.0019$ $0.0020$ $0.0021$ $0.0022$ $0.0027$ $0.0049$
LT Obj Mean $0.778$ $1.175$ $0.608$ $0.932$ $1.627$ $0.729$ $0.254$ $6.651$
Obj Median $0.461$ $0.786$ $0.601$ $0.916$ $1.619$ $0.700$ $0.173$ $6.368$
% Infeasible $0\%$ $1\%$ $0\%$ $0\%$ $3\%$ $0\%$ $0\%$ $5\%$
Time (Sec) $0.0019$ $0.0019$ $0.0019$ $0.0020$ $0.0022$ $0.0022$ $0.0026$ $0.0054$
EX Obj Mean $0.001$ $-0.182$ $-0.453$ $1.649$ $256.926$
Obj Median $-0.095$ $-0.310$ $-0.463$ $-0.052$ $134.620$
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$ $0\%$
% Unsolved $0\%$ $0\%$ $0\%$ $0\%$ $86\%$ $100\%$ $100\%$ $100\%$
Time (Sec) $0.1168$ $2.454$ $994.912$ $1001.52$ $1000.56$
RR Obj Mean $-0.047$ $-0.168$ $-0.464$ $-1.039$ $-2.068$ $-3.990$ $-9.391$
Obj Median $-0.089$ $-0.325$ $-0.476$ $-1.215$ $-2.307$ $-4.327$ $-9.221$
% Infeasible $64\%$ $86\%$ $97\%$ $100\%$ $100\%$ $100\%$ $100\%$
% Unsolved $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $100\%$
Time (Sec) $0.216$ $0.411$ $0.996$ $1.189$ $4.600$ $54.009$ $449.02$
N1 Obj Mean $1.690$ $1073.90$ $42.13e$ 4 $63.72e$ 6 $4411.45$
Obj Median $0.183$ $0.557$ $2.222$ $45.847$ $155.254$
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$ $0\%$
% Unsolved $0\%$ $0\%$ $0\%$ $0\%$ $86\%$ $100\%$ $100\%$ $100\%$
Time (Sec) $0.040$ $0.103$ $0.144$ $8.968$ $940.43$
@@ 6.4 Multi-Dimensional Mixed-Integer Rosenbrock Problem
The high-dimensional mixed-integer Rosenbrock problem is a challenging benchmark adapted from the classic Rosenbrock function, extended with integer variables, non-linear constraints, and parametric variations. It evaluates scalability and the ability to handle complex optimization landscapes. All parameters and the constraint structure, are described in Appendix E.
@@
Table 4: Results for the Mixed-Integer Rosenbrock Problem. The number of decision variables varies from 2 to 20,000, while the number of constraints is 4. See the caption of Table 2 for details.
@@
RC Obj Mean / Median $23.84/21.77$ $53.94/50.48$ $550.95/536.18$ $6034.64/5864.99$ $46.05e4/5.07e$ 4
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$ $31\%$
Time (Sec) $0.0019$ $0.0019$ $0.0022$ $0.0032$ $0.0126$
LT Obj Mean / Median $23.40/21.00$ $54.97/52.34$ $563.67/541.81$ $5465.80/5400.32$ $51.86e5/1.74e$ 5
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$ $45\%$
Time (Sec) $0.0020$ $0.0019$ $0.0022$ $0.0031$ $0.0119$
EX Obj Mean / Median $19.62/18.20$ $64.67/59.16$ $908.818.43e5/$ 908.81 $9262.094.70e10/$ 9262.09 $51.09e15/1.03e$ 5
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$ $0\%$
% Unsolved $0\%$ $0\%$ $0\%$ $4\%$ $22\%$
Time (Sec) $3.509$ $1004.61$ $1002.20$ $1001.69$ $1040.06$
RR Obj Mean / Median $22.24/22.19$ $12036.99/51.17$ $501.901.43e4/$ 501.90 $5436.602.10e6/$ 5436.60 $61.75e8/7.03e$ 6
% Infeasible $45\%$ $41\%$ $18\%$ $1\%$ $4\%$
% Unsolved $0\%$ $0\%$ $42\%$ $93\%$ $78\%$
Time (Sec) $0.181$ $0.557$ $1.240$ $9.233$ $1064.12$
N1 Obj Mean / Median $40.37/27.93$ $87.83/77.34$ $957.423.72e8/$ 957.42 $9379.378.27e12/$ 9379.37 $51.22e15/1.03e$ 5
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$ $0\%$
% Unsolved $0\%$ $0\%$ $0\%$ $5\%$ $22\%$
Time (Sec) $0.032$ $0.081$ $0.261$ $71.911$ $782.052$
@@ 6.6 Effect of Training Sample Size
To assess the impact of sample size, we trained the model on datasets of 800, 8,000, and 80,000 instances with 100 penalty weight, adjusting training epochs to 2000, 200, and 20 (with early stopping) to ensure comparable iterations. All other hyperparameters remained consistent to isolate the effect of sample size.
@@
As shown in Figure 4, increasing the sample size yields significant improvements in both objective values and feasibility. With 80,000 samples for training, the infeasibility ratio was reduced to 5% on the test set, demonstrating better generalization to unseen instances. This emphasizes the critical role of sufficient sample size and demonstrates the scalability advantage of our self-supervised framework.
@@ 7 Conclusion
Our method enables efficient heuristic solutions for large-scale parametric MINLPs, achieving better performance and computational efficiency, though feasibility is not guaranteed. Future work could explore improving feasibility through alternative constraint-handling techniques or post-processing. For certain problem classes, a subset of constraints could be relaxed into the loss function while directly optimizing over the rest using differentiable optimization layers Agrawal et al. (2019). Additionally, redesigning neural network architectures to handle varying instance parameters and decision variables is a promising direction, leveraging set-based, permutation-equivariant architectures such as graph neural networks Cappart et al. (2023); Dumouchelle et al. (2024); Chen et al. (2022b; 2024).
@@ Appendix A Gumbel-Sigmoid Trick
The Gumbel-Sigmoid trick is a stochastic approximation method that enables gradient-based optimization for discrete variables. It introduces controlled random perturbations to the logits $h$, producing a continuous relaxation of the binary decision. The method relies on noise sampled from the Gumbel distribution, which is defined as:
@@
where $U$ is a random variable drawn from the uniform distribution over the interval $[0,1]$. This noise $g$ is added to the logits to introduce randomness while preserving differentiability.
The Gumbel-Sigmoid function provides a soft, differentiable approximation of binary outputs as:
@@
where $g_{1}$ and $g_{2}$ are independent samples from the Gumbel distribution, and $\tau>0$ is the temperature parameter that controls the smoothness of the approximation. For large values of $\tau$, the output of the Gumbel-Sigmoid function is smooth and probabilistic, resembling a sigmoid function. As $tau\to 0$, the output approaches a hard binary decision, mimicking the behavior of a step function.
In optimization contexts, this property allows the Gumbel-Sigmoid trick to approximate discrete variables while maintaining differentiability, facilitating gradient-based methods. The noise introduced by $g_{1}$ and $g_{2}$ promotes exploration, potentially helping to escape poor local minima during training—an issue that is particularly common in discrete optimization problems.
In our experiments, we set $\tau=1$ for simplicity. This setting has been empirically shown to perform well across various learning-to-optimize tasks but can be further tuned to balance the trade-off between exploration and exploitation.